OpenStudy (mony01):

sketch the curve with the given polar equation by first sketching the graph of r as a function of theta in Cartesion coordinates. r=-2sin(theta)

OpenStudy (anonymous):

this is one way to do it: multiply both sides by r where r=/= 0 \[r^2 = 2rsin(\theta)\] use the fact that r^2 = x^2 + y^2 = r^2 and that rsin(theta) = y so your equation is:\[x^2 + y^2 = 2y\]\[x^2 + y^2 - 2y = 0\] you'll want to complete the square to get it in the general form of a circle equation. should i keep going or should i leave you the fun part?

OpenStudy (mony01):

can you please keep going?

OpenStudy (anonymous):

completing the square:\[x^2 + y^2 - 2y = 0\]\[x^2 + [(y - 1)^2 - 1] = 0\]\[x^2 + (y -1)^2 = 1\] the general equation of a circle:\[(x - h)^2 + (y - k)^2 = r^2\]where r is the radius and the center of the circle is the point (h, k) in our case, we have a circle of radius 1 centered at (0, 1). do you know how to you draw it?

OpenStudy (mony01):

not really

OpenStudy (mony01):

wait is r=-2 sin theta

OpenStudy (anonymous):

my drawing was wrong, sorry. it doesn't doesn't change much though -y = -rsin(theta) so we have x^2 + y^2 + 2y = 0 we end up with: x^2 + (y + 1)^2 = 1 this would be centered at (0, -1)

OpenStudy (anonymous):



OpenStudy (anonymous):

i hope this helps. re-reading it might help if it didn't. i have to go make food; starving!

OpenStudy (mony01):

thank you so much